# ROK Defense M&S in the Age of Hyperscale AI: Concepts, Challenges, and Future Directions

### v1

@@ Abstract
Integrating hyperscale AI into national defense modeling and simulation (M&S) is crucial for enhancing strategic and operational capabilities. We explore how hyperscale AI can revolutionize defense M&S by providing unprecedented accuracy, speed, and the ability to simulate complex scenarios. Countries such as the United States and China are at the forefront of adopting these technologies and are experiencing varying degrees of success. Maximizing the potential of hyperscale AI necessitates addressing critical challenges, such as closed networks, long-tail data, complex decision-making, and a shortage of experts. Future directions emphasize the adoption of domestic foundation models, the investment in various GPUs / NPUs, the utilization of big tech services, and the use of open source software. These initiatives will enhance national security, maintain competitive advantages, and promote broader technological and economic progress. With this blueprint, the Republic of Korea can strengthen its defense capabilities and stay ahead of the emerging threats of modern warfare.
Index Terms: hyperscale AI, defense M&S, defense innovation, defense policy
@@ I Introduction
In recent years, integrating AI into defense M&S is crucial as global security threats evolve [1, 2]. The ability to simulate complex scenarios is crucial, and AI can revolutionize defense M&S with its advanced algorithms [3]. These systems provide unprecedented accuracy and speed, leading to better strategic decisions. Improved outcomes ensure greater national security, making it essential to stay ahead of adversaries. AI-driven models precisely predict enemy movements, while accelerated simulations allow real-time decisions [4]. Increased computational power supports complex analyses, while data-driven insights improve tactical responses [5]. In addition, AI can efficiently streamline resource allocation [6].
Countries around the world are recognizing the strategic importance of AI in defense. Nations such as United States (U.S.), China, Japan, European Union, and United Kingdom have comprehensive AI strategies, which highlight the role of AI in modern warfare. The Republic of Korea (ROK) has also announced defense AI strategies to adopt AI to remain competitive. National defense strategies are rapidly evolving, and AI integration is now a global priority. Falling behind can have severe consequences, especially as the pace of AI adoption accelerates. Therefore, early adopters will gain significant advantages, with investment in AI technologies increasing worldwide [7].
Recently, many countries have taken advantage of AI to address previously unexplored areas in traditional defense M&S [8]. This integration not only solves uncharted challenges, but also automates tasks that previously required significant manual effort and time. For example, the ROK Ministry of National Defense (MND) has developed and is using an LLM called GeDAI (Generative Defense AI), as shown in Fig. 1, to streamline various processes. This approach demonstrates a shift towards more efficient, AI-driven solutions that enhance operational capabilities and reduce reliance on human intervention in critical defense tasks. Furthermore, the implementation of AI technologies such as GeDAI is expected to accelerate decision-making processes and improve strategic planning within the defense sectors.
@@
  • Identification of the challenges that need to be addressed for integrating AI into Defense M&S in ROK. These challenges include technical, operational, and policy barriers unique to ROK’s defense landscape.
@@ II-A Defense M&S
Defense M&S is a critical tool that enables military organizations to create virtual representations of real-world systems, processes, and scenarios [9, 10]. By simulating combat situations, training exercises, and equipment performance, defense M&S allows for thorough analysis and preparation without the risks and costs associated with live experiments. As shown in Fig. 2, the hierarchical levels of simulation range from constructive simulations at the base, through virtual simulations in the middle, to live simulations at the apex. For example, using defense M&S to predict the outcomes of battle scenarios can significantly reduce the expenses of conducting full-scale field exercises, as potential issues can be identified and addressed within the simulation environment.
@@ II-B Hyperscale AI
Hyperscale AI operates on an unprecedented scale, utilizing massive computational resources and vast datasets to solve complex problems [11]. As illustrated in Fig. 3, possible applications of hyperscale AI include compter vision, natural language, multimodal, tabular, and audio processing. Hyperscale AI is being used to enhance predictive analytics, automate decision-making processes, and provide deep insights through deep learning and machine learning algorithms. By applying hyperscale AI to defense M&S, it is possible to conduct wargames with reduced human intervention as COA-GPT [12]. This allows simulation of numerous scenarios, strategic decision making, and outcome predictions, thereby decreasing the time and resources required for human-led simulations.
@@
Sep. 2023 Mistral 7B Large Language Model Mistral AI
Oct. 2023 Ernie 4.0 Large Language Model Baidu
@@ III-A Closed Network
One of the major challenges in adopting AI in defense M&S is closed networks. In practice, military secrets are hard to utilize for AI training in a closed network environment. However, open source tools and libraries are foundational for modern AI development, offering cutting-edge algorithms and efficient frameworks that accelerate innovation. Without access to these resources, researchers are compelled to invest time and effort in developing basic functionalities from scratch, leading to redundant work and slower progress. This not only delays project timelines, but also increases costs, as proprietary solutions require continuous maintenance and updates. Furthermore, limited interaction with the open source community isolates defense researchers from collaborative progress, impeding information sharing and the adoption of best practices.
@@ III-B Long-Tail Data
Long-tail data challenges significantly hinder AI development in defense M&S, particularly when training data for rare or adversarial scenarios are limited or unavailable. For example, while extensive data may be available for ROK unmanned aerial vehicles (UAVs), there is a lack of data on enemy UAVs, limiting the AI’s ability to accurately model or predict adversarial behavior. This absence impedes the development of robust AI models equipped to handle a broad spectrum of scenarios, particularly those involving adversarial equipment or tactics that remain insufficiently documented or inaccessible. Consequently, AI systems may underperform in critical situations where understanding enemy actions is vital.
@@ III-D Scarcity of AI Experts
Furthermore, the rotational position system for military officers in ROK exacerbates the shortage of specialized AI personnel. Officers are frequently reassigned to different roles, preventing them from developing deep expertise in the AI field. This lack of continuity and specialization undermines the military’s capacity to cultivate and retain skilled professionals who can drive AI initiatives forward. The absence of dedicated AI specialists means that defense projects may suffer from a lack of technical proficiency, leading to suboptimal outcomes. Addressing this issue is essential for building a robust AI capability within the military, ensuring that personnel can develop and apply specialized knowledge effectively.
@@ IV-A Adopt Domestic Foundation Models
A potential future direction for integrating AI into defense M&S in ROK is the active adoption of foundation models such as HyperCLOVA X and Exaone 2.0 developed by domestic companies, mirroring strategies employed by U.S.. By leveraging the expertise of private sector AI developers, the military can accelerate the implementation of advanced AI technologies. These foundation models, already robust and well-tested, can provide a solid starting point for defense applications, reducing development time and costs. Collaborating with Korean AI companies can also foster innovation and ensure that the models are tailored to the specific needs and contexts of ROK’s defense landscape.
@@ IV-B Involve Major IT Corporations
Mirroring the strategy of the U.S. Defense Innovation Unit (DIU), a key future direction for ROK involves encouraging its leading IT corporations—such as NAVER, Samsung, and SK—to actively participate in R&D initiatives within the defense M&S sector. By opening doors to these industry leaders who are pioneering and dominating the latest IoT market, the military can leverage cutting-edge technologies to enhance defense M&S performance. This collaboration enables the integration of advanced IoT solutions into military simulations and operations, fostering innovation and bridging the gap between civilian technological advancements and defense applications. Such partnerships can accelerate the development of sophisticated simulation models and AI systems tailored to the military’s specific needs.
Engaging large IT companies in defense projects also allows the military to rapidly acquire new weapon systems by adapting existing corporate solutions for military use. These companies possess cutting-edge technologies and specialized expertise that can be optimized to address defense requirements, thereby reducing both the development time and associated costs. By collaborating with IT giants, the military can ensure that its defense IoT infrastructure is built using the most efficient and secure technologies available. This not only enhances operational capabilities, but also addresses the shortage of specialized AI personnel by tapping into the skilled workforce of these corporations. Ultimately, this approach can strengthen national defense while fostering a synergistic relationship between the military and the private sector.
@@ IV-C Diversify GPU/NPU Infrastructure
The need for a diverse GPU/NPU infrastructure is a critical future direction for integrating AI into defense M&S in ROK. Currently, procuring NVIDIA AI hardware is challenging, and over-reliance on a single company’s products can lead to vulnerabilities where national defense capabilities might be influenced or constrained by corporate decisions. This dependency poses risks to security and operational readiness, as any disruptions in the supply chain or changes in corporate policies could adversely affect the military’s AI capabilities. Therefore, it is imperative to reduce reliance on a single vendor to ensure robustness and autonomy in defense technology.
By adopting and utilizing GPUs and NPUs from multiple sources such as Intel, Samsung Electronics, and AMD, the military can build a more resilient and flexible AI infrastructure. This diversification not only mitigates the risks associated with supply chain disruptions but also fosters competitive innovation among suppliers, potentially leading to better performance and cost savings. Incorporating hardware from domestic companies like Samsung Electronics also supports the national industry and can enhance security through closer collaboration. A multivendor approach ensures that the defense sector is not held hostage by any single company’s technology, thereby strengthening national security and technological sovereignty.
@@ IV-D Leverage Open Source Software & Data
To enhance AI-based M&S for confidential data within the military, it is essential to utilize not only software and data from private companies, but also publicly available software and data. By integrating open source tools and datasets, the military can develop robust AI models without the need to export sensitive data, thus maintaining security while benefiting from the latest technological advancements. This approach allows for the internal development of AI solutions tailored to the military’s specific needs, leveraging the collective innovation of the global AI community.
Achieving this integration requires the elimination of complex approval procedures when using externally available software and data internally. By streamlining bureaucratic processes and providing an environment relatively free from punitive measures, researchers and developers can more rapidly adopt new technologies in defense M&S. This reduction in administrative barriers encourages innovation and agility, enabling the military to stay ahead of technological advancements. Creating a more permissive environment for technology adoption not only accelerates development but also attracts talent eager to work in a supportive and forward-thinking environment.
@@ V Conclusion
Integrating hyperscale AI into ROK’s defense M&S offers significant opportunities and challenges. Addressing obstacles such as closed networks, long-tail data scarcity, complex decision-making processes, and a shortage of AI experts is crucial for advancing the nation’s defense capabilities. Through the adoption of domestic foundation models, the participation of major IT corporations in defense R&D, diversification of GPU/NPU infrastructure, and utilization of open source software and data, ROK can significantly enhance its defense M&S capabilities. These strategic directions not only address existing challenges but also promote innovation, enhance operational readiness, and safeguard technological sovereignty. Embracing these initiatives will enable ROK to strengthen national security, stay ahead of emerging threats, and contribute to broader technological and economic growth in the age of hyperscale AI.
@@ VI Acknowledgment
This research was supported by the MSIT (Ministry of Science and ICT), Korea, under the ITRC (Information Technology Research Center) support program (IITP-2024-2020-0-01787) supervised by the IITP (Institute of Information & Communications Technology Planning & Evaluation).

### v2

@@ Abstract
Integrating hyperscale AI into national defense M&S(Modeling and Simulation), under the expanding IoMDT(Internet of Military Defense Things) framework, is crucial for boosting strategic and operational readiness. We examine how IoMDT-driven hyperscale AI can provide high accuracy, speed, and the ability to simulate complex, interconnected battlefield scenarios in defense M&S. Countries like the United States and China are leading the adoption of these technologies, with varying levels of success. However, realizing the full potential of hyperscale AI requires overcoming challenges such as closed networks, sparse or “long-tail” data, complex decision-making processes, and a shortage of experts. Future directions highlight the need to adopt domestic foundation models, expand GPU/NPU investments, leverage large tech services, and employ open source solutions. These efforts will enhance national security, maintain a competitive edge, and spur broader technological and economic growth. With this blueprint, the Republic of Korea can strengthen its defense posture and stay ahead of emerging threats in modern warfare.
Index Terms: IoMDT, hyperscale AI, defense M&S, defense innovation, defense policy
@@ I Introduction
In recent years, integrating AI technologies into defense M&S(Modeling and Simulation) is crucial as global security threats evolve[1, 2]. The ability to simulate complex scenarios is crucial, and AI can revolutionize defense M&S with its advanced algorithms [3]. These systems provide unprecedented accuracy and speed, leading to better strategic decisions. Improved outcomes ensure greater national security, making it essential to stay ahead of adversaries. IoMDT(Internet of Military Defense Things)-equipped AI-driven models precisely predict enemy movements by fusing data from interconnected military IoT assets, while accelerated simulations allow near real-time decisions across heterogeneous platforms [4]. Increased computational power supports complex analyses, while data-driven insights improve tactical responses [5]. In addition, AI can efficiently streamline resource allocation [6].
Countries around the world are recognizing the strategic importance of AI in defense. Nations such as United States (U.S.), China, Japan, European Union, and United Kingdom have comprehensive AI strategies, which highlight the role of AI in modern warfare. Specifically, these nations integrate IoT-based defense architectures to better coordinate troops, optimize supply lines, and automate early warnings. The Republic of Korea (ROK) has also announced defense AI strategies to adopt AI to remain competitive. A key part of these strategies involves leveraging IoMDT to optimize force deployment, logistics, and battlefield command through interconnected devices. National defense strategies continue to advance at an unprecedented pace and AI integration is now a global priority. Delayed implementation of AI technologies can have severe consequences for military capabilities, particularly given the rapid advancement of AI applications in defense systems. Therefore, early adopters will gain significant advantages, with investment in AI technologies increasing worldwide [7].
Recently, many countries have taken advantage of AI and IoMDT technologies to address previously unexplored areas in traditional defense M&S, ranging from soldier health-monitoring wearables to autonomous surveillance systems [8]. This integration not only solves uncharted challenges, but also automates tasks that previously required significant manual effort and time. For example, the ROK Ministry of National Defense (MND) has developed and is using an LLM called GeDAI (Generative Defense AI), as shown in Fig. 1, to streamline various processes. Specifically, GeDAI function like ChatGPT, enabling the ROK military to easily access the information they need. This approach demonstrates a shift towards more efficient, AI-driven solutions that enhance operational capabilities and reduce reliance on human intervention in critical defense tasks. By pairing IoMDT-based data with large-scale generative or predictive models enables real-time, data-driven analysis and rapid field adaptation, moving beyond traditional simulations. Furthermore, the implementation of AI technologies such as GeDAI is expected to strengthen both the efficiency of decision-making processes and the effectiveness of strategic planning within the defense sectors.
@@
  • Identification of the challenges in integrating AI into Defense M&S in ROK. These challenges include technical, operational, and policy characteristics and requirements specific to ROK’s defense environment.
@@ II Concepts
In this section, we introduce the fundamental concepts of defense M&S and hyperscale AI, as well as their relationship within an IoMDT environment.
@@ II-A Defense M&S
Defense M&S is a critical tool that enables military organizations to create virtual representations of real-world systems, processes, and scenarios [9, 10]. By simulating combat situations, training exercises, IoMDT device inter connectivity, and equipment performance, defense M&S allows for thorough analysis and preparation without the risks and costs associated with live experiments. As shown in Fig. 2, the hierarchical levels of simulation range from constructive simulations at the base, through virtual simulations in the middle, to live simulations at the apex. For example, using defense M&S to predict the outcomes of battle scenarios can significantly reduce the expenses of conducting full-scale field exercises, as potential issues can be identified and addressed within the simulation environment.
@@ II-B Hyperscale AI
Hyperscale AI, when paired with IoMDT sensor data, operates on an unprecedented scale, leveraging massive computational resources and diverse battlefield datasets to solve complex operational problems [11]. As illustrated in Fig. 3, possible applications of hyperscale AI include compter vision, natural language, multimodal, tabular, and audio processing. Hyperscale AI is being used to enhance predictive analytics, automate decision-making processes, and provide deep insights through machine learning algorithms. For instance, applying hyperscale AI to defense M&S, it is possible to conduct wargames with reduced human intervention as COA-GPT [12]. This allows simulation of numerous scenarios, strategic decision making, and outcome predictions, thereby decreasing the time and resources required for human-led simulations.
@@
Nov. 2024 Solar Pro Large Language Model Upstage
Dec. 2024 EXAONE 3.5 Large Language Model LG AI Research
@@
Nov. 2023 Claude 2.1 Large Language Model Anthropic
Nov. 2023 Grok-1 Large Language Model xAI
@@
Aug. 2024 Grok-2 Large Language Model xAI
Sep. 2024 LLaMA 3.2 Large Language Model Meta
Dec. 2024 LLaMA 3.3 Large Language Model Meta
Dec. 2024 Gemini 2.0 Large Language Model Google
@@
The ROK military is focusing on four key domains for IoMDT-driven AI applications, as illustrated in Fig. 4. Within these domains, ROK can develop specific AI capabilities tailored to military needs. By leveraging hyperscale AI technologies from domestic tech companies like NAVER, Kakao, and Samsung Electronics, these military domains can be significantly enhanced. The advanced language processing and multimodal capabilities of commercial AI models could improve battlefield analysis and decision-making processes by processing diverse IoMDT data streams in real-time. Such integration would enhance both operational efficiency and combat effectiveness while maintaining security through the military’s specialized IoMDT platforms.
II-C IoMDT for Next-Generation Simulations
While commercial IoT systems focus on consumer-centric devices, IoMDT centers on mission-critical devices, sensors, and applications specialized for military operations. By merging IoMDT data with hyperscale AI capabilities, defense M&S can capture a more comprehensive digital twin of the battlefield. For instance, real-time status from soldier-worn biosensors, UAV imagery, radar signals, and supply chain IoT nodes feed directly into high-fidelity simulations. This integrated architecture offers an adaptive loop: sensor data update simulations, while simulation outcomes inform AI-driven optimization or predictive analytics for resource allocation, mission planning, and threat detection. Moreover, IoMDT frameworks require stringent cybersecurity and reliability measures, distinct from commercial IoT. The robust encryption, secure authentication, and redundant network topologies must ensure that IoMDT data seamlessly reaches the M&S pipeline without compromising battlefield security.
Figure 4: Comprehensive illustration of ROK’s IoMDT architecture.
@@ III Core Challenges
In this section, we identify and discuss the principal obstacles that arise when integrating AI into ROK defense M&S.
@@ III-A Closed Network
One of the major challenges in adopting IoMDT-enhanced AI in defense M&S is operating within closed networks, where the secure integration of interconnected devices is restricted. In practice, military secrets are hard to utilize for AI training in a closed network environment. However, open source tools and libraries are foundational for modern AI development, offering cutting-edge algorithms and efficient frameworks that accelerate innovation. Without access to these resources, researchers are compelled to invest time and effort in developing basic functionalities from scratch, leading to redundant work and slower progress. This not only delays project timelines, but also increases costs, as proprietary solutions require continuous maintenance and updates. Furthermore, limited interaction with the open source community isolates defense researchers from collaborative progress, impeding information sharing and the adoption of best practices.
@@ III-B Long-Tail Data
Long-tail data challenges significantly hinder IoMDT-driven AI development in defense M&S, particularly when sensor inputs and operational logs from rare or adversarial scenarios are limited or unavailable. For example, while extensive data may be available for ROK unmanned aerial vehicles (UAVs), there is a lack of data on enemy UAVs, limiting the AI’s ability to accurately model or predict adversarial behavior. This absence impedes the development of robust AI models equipped to handle a broad spectrum of scenarios, particularly those involving adversarial equipment or tactics that remain insufficiently documented or inaccessible. Consequently, AI systems may underperform in critical situations where understanding enemy actions is vital.
@@ III-D Scarcity of AI Experts
Furthermore, the rotational position system for military officers in ROK exacerbates the shortage of specialized AI personnel. The officers are frequently reassigned to different roles, preventing them from acquiring sufficient expertise in the AI field. This lack of continuity and specialization undermines the military’s capacity to cultivate and retain skilled professionals who can drive AI initiatives forward. The absence of dedicated AI specialists means that defense projects may suffer from a lack of technical proficiency, leading to suboptimal outcomes. Addressing this issue is essential for building a robust AI capability within the military, ensuring that personnel can develop and apply specialized knowledge effectively.
@@ IV Future Directions
In this section, we propose potential solutions and directions to achieve hyperscale AI-based defense M&S in ROK as shown in Fig. 5.
@@ IV-A Adopt Domestic Foundation Models
A potential future direction for integrating AI into IoMDT-based defense M&S in ROK is the active adoption of foundation models such as HyperCLOVA X and EXAONE 3.5 developed by domestic companies, mirroring strategies employed by U.S. DoD. By leveraging the expertise of private sector AI developers, the military can accelerate the implementation of advanced AI technologies. These foundation models, already robust and well-tested, can provide a solid starting point for defense applications, reducing development time and costs. Collaborating with Korean AI companies can also foster innovation and ensure that the models are tailored to the specific needs and contexts of ROK’s defense landscape.
@@ IV-B Involve Major IT Corporations
Mirroring the strategy of the U.S. Defense Innovation Unit (DIU), a key future direction for ROK involves encouraging its leading IT corporations—such as NAVER, Samsung Electronics, and SKT to actively participate in R&D initiatives within the defense M&S sector. Through strategic collaboration with these industry leaders who are pioneering the IoT and IoMDT markets, the military can leverage cutting-edge connectivity, sensor fusion, and cloud-edge AI technologies to enhance defense M&S performance. This collaboration enables the integration of advanced IoT solutions into wargame and operations, fostering innovation and bridging the gap between civilian technological advancements and defense applications. Such partnerships can accelerate the development of sophisticated simulation models and AI systems tailored to the military’s specific needs.
Engaging large IT companies in defense projects also allows the military to rapidly acquire new weapon systems by adapting existing corporate solutions for military use. These companies possess cutting-edge technologies and specialized expertise that can be optimized to address defense requirements, thereby reducing both the development time and associated costs. By collaborating with leading technology enterprises in the IoMDT sector, the military can ensure that its defense infrastructure is built using the most efficient and secure technologies available. This not only enhances operational capabilities, but also addresses the shortage of specialized AI personnel by tapping into the skilled workforce of these corporations. Ultimately, this approach can strengthen national defense while fostering a synergistic relationship between the military and the private sector.
@@ IV-C Diversify GPU/NPU Infrastructure
The need for a diverse GPU/NPU infrastructure is a critical future direction for integrating AI into IoMDT-based defense M&S in ROK, ensuring robust on-site computing for the vast sensor networks deployed in modern battle environments. Currently, getting NVIDIA AI hardware is challenging, and over-reliance on a single company’s products can lead to vulnerabilities where national defense capabilities might be influenced or constrained by corporate decisions. This dependency poses risks to security and operational readiness, as any disruptions in the supply chain or changes in corporate policies could adversely affect the military’s AI capabilities. Therefore, it is imperative to reduce reliance on a single vendor to ensure robustness and autonomy in defense technology.
By adopting and utilizing GPUs and NPUs from multiple sources such as Gaudi series of Intel, the military can build a more resilient and flexible AI infrastructure. This diversification not only mitigates the risks associated with supply chain disruptions but also fosters competitive innovation among suppliers, potentially leading to better performance and cost savings. Incorporating hardware from domestic companies like Samsung Electronics also supports the national industry and can enhance security through closer collaboration. A multivendor approach ensures that the defense sector is not held hostage by any single company’s technology, thereby strengthening national security and technological sovereignty.
@@ IV-D Leverage Open Source Software & Data
Enhancing AI-based M&S for confidential military data requires leveraging both proprietary and open source software and data. Integrating open source tools and datasets enables the development of robust AI models without compromising sensitive data, ensuring security while leveraging cutting-edge innovations. This strategy facilitates the in-house development of AI solutions customized for military requirements, harnessing the worldwide AI community’s collaborative progress.
Achieving this integration requires the elimination of complex approval procedures when using externally available software and data internally. Simplifying administrative procedures and fostering a supportive environment accelerates the adoption of new technologies in defense M&S. Reducing administrative hurdles promotes innovation and agility, keeping the military at the forefront of technological progress. Cultivating a conducive atmosphere for technology adoption expedites development and draws talented individuals seeking a progressive and encouraging workplace.
@@ V Conclusion
Integrating hyperscale AI and IoMDT into ROK’s defense M&S offers significant opportunities and challenges, laying the groundwork for autonomous operations, enhanced situational awareness, and robust tactical decision support. Addressing obstacles such as closed networks, long-tail data scarcity, complex decision-making processes, and a shortage of AI experts is crucial for advancing the nation’s defense capabilities. Through the adoption of domestic foundation models, the participation of major IT corporations in defense R&D, diversification of GPU/NPU infrastructure, and utilization of open source software and data, ROK can significantly enhance its defense M&S capabilities. These strategic directions address existing challenges while promoting innovation and operational readiness. This approach will enable ROK to strengthen national security and contribute to broader technological and economic growth in the age of hyperscale AI.
@@ VI Acknowledgment
This research was supported by the MSIT (Ministry of Science and ICT), Korea, under the ITRC (Information Technology Research Center) support program (IITP-2025-2020-0-01787) supervised by the IITP (Institute of Information & Communications Technology Planning & Evaluation). Also, this research was supported in part by the NAVER-Intel Co-Lab. The work was conducted by KAIST and reviewed by both NAVER and Intel.